OpenStudy (anonymous):

Why does (2-sin^2(3θ))/(sin^2(3θ)) = 2/(sin^2(3θ)) -1

OpenStudy (zehanz):

Because you can split up the fraction in this way:\[\frac{ a-b }{ b }=\frac{ a }{ b }-\frac{ b }{ b }=\frac{ a }{ b }-1\]

OpenStudy (anonymous):

Ok, thanks, just another dugh moment in life... lol

OpenStudy (zehanz):

yw!